# A determinantal formula for orthosymplectic Schur functions

### v1

@@ 2. Preliminaries
$\mathrm{sp}^{(n-1,1)}(x_{1},\dots,x_{n-1}|x_{n})=\sum_{T\in\operatorname{SPT}^{(n-1,1)}(\lambda)}\operatorname{wt}(T).$
@@
For a partition of length at most $n$, let $A_{\lambda}=(a_{ij})$ be a $n\times n$ matrix such that
@@
Comparing the determinantal formulas in 1.4 and 2.10, we get the following result.
@@
$\mathrm{sp}^{(n-1,1)}_{\lambda}(x_{1},\dots,x_{n-1}|x_{n})=\operatorname{spo}_{\lambda}(x_{1},\dots,x_{n-1},x_{n}^{-1}/-x_{n}^{-1}).$
@@ 3. Determinant formula for orthosymplectic Schur functions
In this section, we give a proof of 1.2. We first prove an equivalent statement of 1.2 given in 3.3. Before that, let us recall the following lemmas.
@@
$\left(\begin{array}[]{cccc}\overline{H}_{l-n}&\overline{H}_{l-n+1}+\overline{H}_{l-n-1}&\dots&\overline{H}_{l-1}+\overline{H}_{l-2n+1}\end{array}\right)\left(\begin{array}[]{c}(-1)^{q-p}e_{q-p}\end{array}\right)\left(\begin{array}[]{c}\zeta_{j}(n)\\ \zeta_{j}(n-1)\\ \vdots\\ \zeta_{j}(1)\end{array}\right),$
where $\overline{H}_{r}=h_{r}(X,\bar{X})$ for all $r\in\mathbb{Z}$.
@@
Let $\lambda$ be a partition such that $\ell(\lambda)\leq n_{1}$ and $\lambda_{1}\leq n_{2}$. Then the $n_{1}+n_{2}$ numbers
$\lambda_{i}+n_{1}-i\quad(1\leq i\leq n_{1}),\qquad\qquad n_{1}-1+j-\lambda_{j}^{\prime}\quad(1\leq j\leq n_{2})$
are a permutation of $\{0,1,\dots,n_{1}+n_{2}-1\}$.
@@
Let $\lambda$ be a partition of length at most $n$ and $k=\min\{j|\lambda_{j}+n+1-j\leq m\}.$ Then the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X/Y)$ is given by
(3.1) $\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{mn-n+k-1}}{D}\det\left(\begin{array}[]{c|c}R&X_{\lambda}\\[-8.5359pt] \\ \hline\cr\\[-8.5359pt] Y_{\lambda}&0\end{array}\right),$
@@
By Jacobi–Trudi type identity for orthosymplectic Schur functions (2.13), we have
(3.2) $\operatorname{spo}_{\lambda}(X/Y)=\det\left(J_{\lambda_{i}-i+j}+J_{\lambda_{i}-i-j+2}\right)_{1\leq i,j\leq n},$
where $J_{r}=\displaystyle\sum_{p=0}^{r}h_{p}(X,\bar{X})e_{r-p}(Y)=\displaystyle\sum_{p=0}^{r}h_{r-p}(X,\bar{X})e_{p}(Y)$ for all $r\geq 0$. Otherwise $J_{r}=0$. Since $e_{p}(Y)=0$ for $p>m$ and $h_{r-p}(X,\bar{X})=0$, for $p>r$,
(3.3) $J_{r}=\displaystyle\sum_{p=0}^{m}h_{r-p}(X,\bar{X})e_{p}(Y).,\quad r\in\mathbb{Z}.$
Substituting the expression for $J_{r}$ from (3.3) in (3.2), we have
$\operatorname{spo}_{\lambda}(X/Y)=\det\left(\sum_{p=0}^{m}(\overline{H}_{\lambda_{i}-i+j-p}+\overline{H}_{\lambda_{i}-i-j+2-p})e_{p}(Y)\right)_{n\times n},$
where $\overline{H}_{r}=h_{r}(X,\bar{X})$ for all $r\in\mathbb{Z}$. Let $k=\text{min}\{j|\lambda_{j}+n+1-j\leq m\}.$ Now applying 3.1, the required orthosymplectic Schur function is given by
$\frac{1}{D}\det\left(\begin{array}[]{c}\left(\displaystyle\sum_{p=0}^{m}\left(x_{j}^{\lambda_{i}+n-i+1-p}-\bar{x}_{j}^{\lambda_{i}+n-i+1-p}\right)e_{p}(Y)\right)_{1\leq i\leq k-1}\\[-8.5359pt] \\ \hline\cr\\[-8.5359pt] \left(\displaystyle\sum_{p=0}^{\lambda_{i}+n-i}\left(x_{j}^{\lambda_{i}+n-i+1-p}-\bar{x}_{j}^{\lambda_{i}+n-i+1-p}\right)e_{p}(Y)\right)_{k\leq i\leq n}\end{array}\right)_{1\leq j\leq n},$
@@
$D=\displaystyle\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j}).$
@@
Consider the permutation $\sigma=(k,\dots,n,1,\dots,k-1)$, in one line notation in $S_{n}$. Then, the number of inversions of $\sigma$ is $(k-1)(n-k+1)$ and it can be seen that
(3.5) $\operatorname{spo}_{\lambda}(X/Y)=\frac{\operatorname{sgn}(\sigma)}{D}\det\left(\begin{array}[]{c|c}(A_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ k\leq j\leq n\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\end{array}\right),$
@@
where the last equality uses the relations given in (3.4). Let $\Lambda\coloneqq\{\lambda_{j}+n-j+1:k\leq j\leq n\}$ and $\Lambda^{\prime}\coloneqq\{n+i-\lambda_{i}^{\prime}:1\leq i\leq m-n+k-1\}$. Note that $\{\Lambda,\Lambda^{\prime}\}$ is a set partition of $\{1,2,\dots,m\}$ by 3.2. So, we can rewrite (3.5) in the following manner.
$\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{b}}{D}\det\left(\begin{array}[]{c|c|c}((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j}))_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&(A_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ k\leq j\leq n\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&\\ \hline\cr&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&{\text{\huge 0}}&{\text{\huge 0}}\end{array}\right),$
@@
So, the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X/Y)$ is given by
@@
where $\phi_{i}(j)=x_{i}^{j}-\bar{x}_{i}^{j}$. Applying $C_{\lambda_{n}+1}\rightarrow C_{\lambda_{n}+1}-\sum_{p=1}^{\lambda_{n}}(-1)^{m-p}e_{\lambda_{n}+1-p}(Y)C_{p}$ and using
@@
the function in (3.6) is the same as
$\frac{(-1)^{b}}{D}\\ \times\left|\begin{array}[]{c|c|c|c}\left((-1)^{m-j}\phi_{i}(j)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&\left(\displaystyle\sum_{p=0}^{j-1}\psi_{i}(p)e_{p}(Y)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda\end{subarray}}&(\phi_{i}(j))_{\begin{subarray}{c}1\leq i\leq n\\ j=\lambda_{n}+1\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&&\\ \hline\cr&&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&{\text{\huge 0}}&((-1)^{m-j}h_{-i+j}(Y))_{\begin{subarray}{c}i\in\Lambda^{\prime}\\ j=\lambda_{n}+1\end{subarray}}&{\text{\huge 0}}\end{array}\right|,$
where $\psi_{i}(p)=\phi_{i}(j-p)$. Successively applying $C_{j}\rightarrow C_{j}-\sum_{p=1}^{j-1}(-1)^{m-p}e_{j-p}C_{p}$, for $j\in\Lambda$, we have
$\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{b}}{D}\det\left(\begin{array}[]{c|c|c}\left((-1)^{m-j}\phi(j)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&(\phi(j))_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&\\ \hline\cr&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&((-1)^{m-j}h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda\end{subarray}}&{\text{\huge 0}}\end{array}\right).$
Rearranging the columns, we have
$\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c}\left((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j})\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\[17.50002pt] \hline\cr\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(Y)\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
@@
Now we prove that the expressions for $\operatorname{spo}_{\lambda}(X/Y)$ in 1.2 and 3.3 are the same.
@@
Assume that (1.1) holds true. Let $a=mn-n+k-1$. Then the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X/Y)$ is given by
(3.7) $\frac{(-1)^{a}}{D_{s}}\det\left(\begin{array}[]{c|c}\left(\displaystyle\frac{x_{i}}{(x_{i}+y_{j})\prod_{q=1}^{m}(\bar{x}_{i}+y_{q})}-\displaystyle\frac{\bar{x}_{i}}{(\bar{x}_{i}+y_{j})\prod_{q=1}^{m}(x_{i}+y_{q})}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&X_{\lambda}\\[17.50002pt] \hline\cr\\ \left(y_{j}^{\lambda^{\prime}_{i}+m-n-i}\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
@@
Applying $R_{i}\rightarrow\prod_{q=1}^{m}(x_{i}+y_{q})(\bar{x}_{i}+y_{q})R_{i}$, for $i\in[1,n]$ and cancelling the term $\prod_{i=1}^{n}\prod_{j=1}^{m}(x_{i}+y_{j})(\bar{x}_{i}+y_{j})$ from the denominator of $D_{s}$, (3.7) can be written as
@@
$Z_{\lambda}=\left({x_{i}^{\lambda_{j}+n-m-j+1}}{\prod_{q=1}^{m}(x_{i}+y_{q})}-{\bar{x}_{i}^{\lambda_{j}+n-m-j+1}}{\prod_{q=1}^{m}(\bar{x}_{i}+y_{q})}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}},$
@@
Applying $C_{j}\rightarrow C_{j}-C_{m}$, $j\in[1,m-1]$ and then using $y_{j}^{p}-y_{m}^{p}=(y_{j}-y_{m})h_{p-1}(y_{j},y_{m})$, the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X/Y)$ is given by

### v2

@@ 2. Preliminaries
$\mathrm{sp}^{(n-1,1)}_{\lambda}(x_{1},\dots,x_{n-1}|x_{n})=\sum_{T\in\operatorname{SP}^{(n-1,1)}(\lambda)}\operatorname{wt}(T).$
Example 2.10.
Suppose $n=2$ and $\lambda=(2,1)$. Then we have the folowing $(1,1)$-symplectic tableaux of shape $\lambda$:
[…]
Therefore,
$\mathrm{sp}^{(1,1)}_{(2,1)}(x_{1}|x_{2})=x_{1}^{2}x_{2}+x_{2}+x_{1}x_{2}^{2}+\bar{x}_{1}^{2}x_{2}+\bar{x}_{1}x_{2}^{2}.$
@@
For a partition of length at most $n$, let $A_{\lambda}=(a_{ij})$ be an $n\times n$ matrix such that
@@
Comparing the determinantal formulas in 1.4 and 2.11, we get the following result.
@@
$\mathrm{sp}^{(n-1,1)}_{\lambda}(x_{1},\dots,x_{n-1}|x_{n})=\operatorname{spo}_{\lambda}(x_{1},\dots,x_{n-1},x_{n}^{-1};-x_{n}^{-1}).$
@@ 3. Determinant formula for orthosymplectic Schur functions
In this section, we give a proof of 1.2. We first prove 3.3, which is equivalent to 1.2. Before that, let us recall the following lemmas.
@@
$\left(\begin{array}[]{c|c}\overline{H}_{l-n}&\left(\overline{H}_{l-n-1+j}+\overline{H}_{l-n+1-j}\right)_{2\leq j\leq n}\end{array}\right)\Big{(}(-1)^{v-u}e_{v-u}\Big{)}_{1\leq u,v\leq n}\Big{(}\zeta_{j}(n+1-u)\Big{)}_{1\leq u\leq n},$
where $\overline{H}_{r}=h_{r}(X,\overline{X})$ for all $r\in\mathbb{Z}$.
@@
Let $\lambda$ be a partition such that $\ell(\lambda)\leq n_{1}$ and $\lambda_{1}\leq n_{2}$. Then $\{\lambda_{i}+n_{1}-i\mid 1\leq i\leq n_{1}\}$ and $\{n_{1}-1+j-\lambda_{j}^{\prime}\mid 1\leq j\leq n_{2}\}$ form a set partition of $\{0,1,\dots,n_{1}+n_{2}-1\}$.
@@
Let $\lambda$ be a partition of length at most $n$ and $k=\min\{j\mid\lambda_{j}+n+1-j\leq m\}.$ Then the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X;Y)$ is given by
(3.1) $\operatorname{spo}_{\lambda}(X;Y)=\frac{(-1)^{mn-n+k-1}}{D}\det\left(\begin{array}[]{c|c}R&X_{\lambda}\\[-8.5359pt] \\ \hline\cr\\[-8.5359pt] Y_{\lambda}&0\end{array}\right),$
@@
By the Jacobi–Trudi type identity for orthosymplectic Schur functions (2.13), we have
(3.2) $\operatorname{spo}_{\lambda}(X;Y)={\det\left(\begin{array}[]{c|c}\left(J_{\lambda_{i}-i+1}\right)_{1\leq i\leq n}&(J_{\lambda_{i}-i+j}+J_{\lambda_{i}-i-j+2})_{\begin{subarray}{c}1\leq i\leq n\\ 2\leq j\leq n\end{subarray}}\end{array}\right)},$
where $J_{r}=\displaystyle\sum_{p=0}^{r}h_{p}(X,\overline{X})e_{r-p}(Y)=\displaystyle\sum_{p=0}^{r}h_{r-p}(X,\overline{X})e_{p}(Y)$ for all $r\geq 0$ and $J_{r}=0$ otherwise. Since $e_{p}(Y)=0$ for $p>m$ and $h_{r-p}(X,\overline{X})=0$, for $p>r$,
(3.3) $J_{r}=\displaystyle\sum_{p=0}^{m}h_{r-p}(X,\overline{X})e_{p}(Y),\quad r\in\mathbb{Z}.$
Substituting the expression for $J_{r}$ from (3.3) in (3.2), the required orthosymplectic Schur function is given by
$\det\left(\begin{array}[]{c|c}\left(\displaystyle\sum_{p=0}^{m}\overline{H}_{\lambda_{i}-i+1-p}e_{p}(Y)\right)_{1\leq i\leq n}&\left(\displaystyle\sum_{p=0}^{m}(\overline{H}_{\lambda_{i}-i+j-p}+\overline{H}_{\lambda_{i}-i-j+2-p})e_{p}(Y)\right)_{\begin{subarray}{c}1\leq i\leq n\\ 2\leq j\leq n\end{subarray}}\end{array}\right)$
where $\overline{H}_{r}=h_{r}(X,\overline{X})$ for all $r\in\mathbb{Z}$. Observe that
(3.4) $\operatorname{spo}_{\lambda}(X;Y)=\det\left(\sum_{p=0}^{m}e_{p}(Y)M^{p}\right),$
@@
$M^{p}=\left(\begin{array}[]{c|c}\left(\displaystyle\overline{H}_{\lambda_{i}-i+1-p}\right)_{1\leq i\leq n}&\left(\displaystyle\overline{H}_{\lambda_{i}-i+j-p}+\overline{H}_{\lambda_{i}-i-j+2-p}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 2\leq j\leq n\end{subarray}}\end{array}\right).$
If $\lambda_{i}+n-i+1-p\geq 0$, then by 3.1, we have
$(M^{p}_{i,v})_{1\leq v\leq n}\left((-1)^{v-u}e_{v-u}\right)_{1\leq u,v\leq n}\left(x_{j}^{n+1-u}-\bar{x}_{j}^{n+1-u}\right)_{1\leq u\leq n}=x_{j}^{\lambda_{i}+n-i+1-p}-\bar{x}_{j}^{\lambda_{i}+n-i+1-p}$
and $M^{p}_{i,v}=0$ for all $1\leq v\leq n$ otherwise. As a consequence, we have
(3.5) $\begin{split}\left(\sum_{p=0}^{m}e_{p}(Y)M^{p}\right)&\Big{(}(-1)^{v-u}e_{v-u}\Big{)}_{1\leq u,v\leq n}\Big{(}x_{j}^{n+1-u}-\bar{x}_{j}^{n+1-u}\Big{)}_{1\leq u\leq n}\\ =&\left(\begin{array}[]{c}\left(\displaystyle\sum_{p=0}^{m}\left(x_{j}^{\lambda_{i}+n-i+1-p}-\bar{x}_{j}^{\lambda_{i}+n-i+1-p}\right)e_{p}(Y)\right)_{1\leq i\leq k-1}\\[-8.5359pt] \\ \hline\cr\\[-8.5359pt] \left(\displaystyle\sum_{p=0}^{\lambda_{i}+n-i}\left(x_{j}^{\lambda_{i}+n-i+1-p}-\bar{x}_{j}^{\lambda_{i}+n-i+1-p}\right)e_{p}(Y)\right)_{k\leq i\leq n}\end{array}\right)_{n\times 1}.\end{split}$
Therefore, using the expressions given in (3.4),(3.5) and the fact that
$\det.\left((-1)^{v-u}e_{v-u}\right)_{1\leq u,v\leq n}=1,$
the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X;Y)$ is given by
(3.6) $\frac{1}{D}\det\left(\begin{array}[]{c}\left(\displaystyle\sum_{p=0}^{m}\left(x_{j}^{\lambda_{i}+n-i+1-p}-\bar{x}_{j}^{\lambda_{i}+n-i+1-p}\right)e_{p}(Y)\right)_{1\leq i\leq k-1}\\[-8.5359pt] \\ \hline\cr\\[-8.5359pt] \left(\displaystyle\sum_{p=0}^{\lambda_{i}+n-i}\left(x_{j}^{\lambda_{i}+n-i+1-p}-\bar{x}_{j}^{\lambda_{i}+n-i+1-p}\right)e_{p}(Y)\right)_{k\leq i\leq n}\end{array}\right)_{1\leq j\leq n},$
where
$D=\det\Big{(}x_{j}^{n+1-u}-\bar{x}_{j}^{n+1-u}\Big{)}_{1\leq u,j\leq n}=\displaystyle\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j}).$
@@
Consider the permutation $\sigma=(k,\dots,n,1,\dots,k-1)\in$ $S_{n}$. The number of inversions of $\sigma$ is $(k-1)(n-k+1)$. Permuting the rows of the determinant in (3.6) by $\sigma$ and then transposing the determinant, we have
(3.8) $\operatorname{spo}_{\lambda}(X;Y)=\frac{\operatorname{sgn}(\sigma)}{D}\det\left(\begin{array}[]{c|c}(A_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ k\leq j\leq n\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\end{array}\right),$
@@
where the last equality uses the relations given in (3.7). Let $\Lambda\coloneqq\{\lambda_{j}+n-j+1\mid k\leq j\leq n\}$, $\Lambda^{\prime}\coloneqq\{n+i-\lambda_{i}^{\prime}\mid 1\leq i\leq m-n+k-1\}$ and
$M\coloneqq\left(\begin{array}[]{c|c|c}((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j}))_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&(A_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ k\leq j\leq n\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&\\ \hline\cr&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&{\text{\huge 0}}&{\text{\huge 0}}\end{array}\right).$
Since
$\det(h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}=1,$
$\det M=(-1)^{n(m-n+k-1)}\det\left(\begin{array}[]{c|c}(A_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ k\leq j\leq n\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\end{array}\right).$
So, we can rewrite (3.8) in the following manner:
$\operatorname{spo}_{\lambda}(X;Y)=\frac{(-1)^{b}}{D}\det\left(\begin{array}[]{c|c|c}((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j}))_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&(A_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ k\leq j\leq n\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&\\ \hline\cr&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&{\text{\huge 0}}&{\text{\huge 0}}\end{array}\right),$
@@
So, the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X;Y)$ is given by
@@
where $\phi_{i}(j)=x_{i}^{j}-\bar{x}_{i}^{j}$. Note that $\{\Lambda,\Lambda^{\prime}\}$ is a set partition of $\{1,2,\dots,m\}$ by 3.2. Applying the elementary column transformation $C_{\lambda_{n}+1}\rightarrow C_{\lambda_{n}+1}-\sum_{p=1}^{\lambda_{n}}(-1)^{m-p}e_{\lambda_{n}+1-p}(Y)C_{p}$ and using
@@
the expression in (3.9) is the same as
${\footnotesize\frac{(-1)^{b}}{D}\left|\begin{array}[]{c|c|c|c}\left((-1)^{m-j}\phi_{i}(j)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&\left(\displaystyle\sum_{p=0}^{j-1}\phi_{i}(j-p)e_{p}(Y)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda\\ j\neq\lambda_{n}+1\end{subarray}}&(\phi_{i}(j))_{\begin{subarray}{c}1\leq i\leq n\\ j=\lambda_{n}+1\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&&\\ \hline\cr&&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&{\text{\huge 0}}&\left((-1)^{m-j}h_{-i+j}(Y)\right)_{\begin{subarray}{c}i\in\Lambda^{\prime}\\ j=\lambda_{n}+1\end{subarray}}&{\text{\huge 0}}\end{array}\right|.}$
Successively applying the elementary column transformations $C_{j}\rightarrow C_{j}-\sum_{p=1}^{j-1}(-1)^{m-p}e_{j-p}C_{p}$, for $j\in\Lambda$, we have
$\operatorname{spo}_{\lambda}(X;Y)=\frac{(-1)^{b}}{D}\det\left(\begin{array}[]{c|c|c}\left((-1)^{m-j}\phi(j)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&(\phi(j))_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&\\ \hline\cr&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&((-1)^{m-j}h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda\end{subarray}}&{\text{\huge 0}}\end{array}\right).$
Taking $(-1)^{m-j}$ out from the $j^{\text{th}}$ column for each $j\in\Lambda$, the orthosymplectic Schur function is given by
$\operatorname{spo}_{\lambda}(X;Y)=\frac{(-1)^{b^{\prime}}}{D}\det\left(\begin{array}[]{c|c|c}\left((-1)^{m-j}\phi(j)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&((-1)^{m-j}\phi(j))_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&\\ \hline\cr&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&(h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda\end{subarray}}&{\text{\huge 0}}\end{array}\right).$
where $b^{\prime}=b+\sum_{j\in\Lambda}(m-j)$. Rearranging the columns, we have
$\operatorname{spo}_{\lambda}(X;Y)=\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c}\left((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j})\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\[17.50002pt] \hline\cr\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(Y)\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
@@
Now we prove that the expressions for $\operatorname{spo}_{\lambda}(X;Y)$ in 1.2 and 3.3 are the same.
@@
Let $a=mn-n+k-1$. Consider the following expression given in (1.1).
${A}\coloneqq\frac{(-1)^{a}}{D_{s}}\det\left(\begin{array}[]{c|c}\left(\displaystyle\frac{x_{i}}{(x_{i}+y_{j})\prod_{q=1}^{m}(\bar{x}_{i}+y_{q})}-\displaystyle\frac{\bar{x}_{i}}{(\bar{x}_{i}+y_{j})\prod_{q=1}^{m}(x_{i}+y_{q})}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&X_{\lambda}\\[17.50002pt] \hline\cr\\ \left(y_{j}^{\lambda^{\prime}_{i}+m-n-i}\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
@@
Applying the elementary row transformations $R_{i}\rightarrow\prod_{q=1}^{m}(x_{i}+y_{q})(\bar{x}_{i}+y_{q})R_{i}$, for $i\in[1,n]$ and cancelling the term $\prod_{i=1}^{n}\prod_{j=1}^{m}(x_{i}+y_{j})(\bar{x}_{i}+y_{j})$ from the denominator of $D_{s}$, the expression for $A$ can be written as
@@
$Z_{\lambda}=\left({x_{i}^{\lambda_{j}+n-m-j+1}}{\prod_{q=1}^{m}(x_{i}+y_{q})}-{\bar{x}_{i}^{\lambda_{j}+n-m-j+1}}{\prod_{q=1}^{m}(\bar{x}_{i}+y_{q})}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}.$
@@
Applying the elementary column transformations $C_{j}\rightarrow C_{j}-C_{m}$, $j\in[1,m-1]$ and then using $y_{j}^{p}-y_{m}^{p}=(y_{j}-y_{m})h_{p-1}(y_{j},y_{m})$ to factor out the term $y_{j}-y_{m}$ for $j\in[1,m-1]$, $A$ is given by